CH/Agen/92/04

Agenda

Charissa Group Meeting

To be held on: 24 August 1992, 11.00 am To be held at: Birmingham, Room E118

Distribution: WNC,NMC,RAC,BRF,GJG,RAH,JSL,VP,WDMR,GT,DLW

1.  Minutes of Meeting of 29 April 1992
2.  Matters Arising/ Actions
3.  Update on NSC/SERC matters  (DLW)
4.  Summary of Grant Spend  (BRF,WDMR,WNC,DLW)
5.  Progress on scintillators  (NMC)
6.  Progress on silicon detectors  (BRF)
7.  Progress on gas supply  (RAC)
8.  Progress on chamber & detector assemblies  (JSL)
9.  Progress on electronics & data acquisition  (RAH)
10.  Developments on hybrid detectors  (RAC)
11.  Developments on event manager & GEC emulator  (RAH,VP)
12.  Update on Recent Analysis  (BRF,WDMR,WNC,DLW)
13.  Discussion of Future Plans
14.  Date for Physics Meeting
15.  Any other business
16.  Date of next meeting

W.N. Catford (Secretary)